# Kalman Filter Enhanced GRPO for Reinforcement Learning-Based Language Model Reasoning

### v3

@@ Abstract
The advantage function is a central concept in RL that helps reduce variance in policy gradient estimates. Recently, for language modeling, Group Relative Policy Optimization (GRPO) was proposed to compute the advantage for each output by subtracting the mean reward, as the baseline, for all outputs in the group. However, it can lead to high variance when the reward advantage is inaccurately predicted. In this work, we propose Kalman Filter Enhanced Group Relative Policy Optimization (KRPO) model, by using lightweight Kalman filtering to dynamically estimate the latent reward baseline and uncertainty. This filtering technique replaces the naive group mean, enabling more adaptive advantage normalization. Our method does not require additional learned parameters over GRPO. This approach offers a simple yet effective way to incorporate multiple outputs of GRPO into advantage estimation, improving policy optimization in settings where highly dynamic reward signals are difficult to model for language models. Through the accuracies and rewards obtained from math question answering and reasoning, we show that using a more adaptive advantage estimation model, KRPO can improve the stability and performance of GRPO. The code is available at https://github.com/billhhh/KRPO_LLMs_RL.
@@ 2.2 Kalman-Filtered Advantage Estimation
We propose replacing the naive group mean baseline with a Kalman filter that maintains an online estimate of the expected reward and its variance to obtain a more accurate estimate of the advantages. Each observed return $r_{i}$ is treated as a noisy measurement of the latent expected reward signal $\hat{x}_{i}$, which we cannot obtain directly.
@@ 2.3 KRPO Algorithm
The KRPO procedure is fully differentiable and can be implemented efficiently upon the GRPO frameworks. After getting the group rewards, instead of receiving the mean value from the group, we fetch the Kalman filtered baseline and estimated uncertainty $P_{i|i}$. By taking into account the posterior variance estimation $P_{i|i}$, the advantage can be more adaptively modeled.
@@
Hyperparameters: KL weight $\beta$, group size $n$, batch size $B$, learning rate $\eta$, clipping $\epsilon$, process noise $Q$, measurement noise $R$ Initialize replay buffer $\mathcal{B}$, optimizer $\text{Adam}(\theta,\eta)$ ;
@@
     Use a Kalman Filter to estimate reward mean $\hat{x}_{i|i}$ and variance $P_{i|i}$ from $\{r_{i}\}_{i=1}^{n}$ ;
@@
Worth noting, in the model training, the observations are completely randomly sampled in each iteration. Moreover, the order of samples in our implementation is arbitrary (e.g., they can be shuffled before processing), so no unintended temporal dependence is imposed on the data. It simply provides a recursive estimator with an accompanying measure of uncertainty. The ordering will thus not effect policy optimization.
Besides, each Kalman filter operation is applied to groups of rollouts. In the KRPO setting, although the reward values for individual samples are discrete, the grouped reward observations, which aggregate multiple rollouts, are not sparse and are much more continuous in distribution. According to the Central Limit Theorem (CLT, 2008), the sum of i.i.d. sampled rewards (with finite mean and finite variance) tends to follow a Gaussian distribution. This is consistent with the assumptions of our KRPO setting.
3 Connections of Group-mean and Kalman Filtering
$\bar{r}=\frac{1}{n}\sum_{i=1}^{n}r_{i}.$ (7)
4 Experiments
4.1 Implementation Details
We adopt four datasets in our experiments, including Arithmetic (Open-Thought, 2024), OpenMathInstruct (Toshniwal et al., 2024), MATH500 (Lightman et al., 2023), and AIME (Veeraboina, 2024). We keep the standard training/testing 80%/20% split ratio for each of the four datasets without model selection. Arithmetic and OpenMathInstruct are divided into three sub-datasets individually: easy, normal, and hard. More details about the datasets used in the paper can be found in the appendix. In the paper, we focus on math tasks. These tasks are suitable for experiments as they require the model to have adequate reasoning ability for resolving difficult math questions (even not easy for humans), and language model reasoning is the main focus of our KRPO paper. Moreover, math tasks have well-defined answers for reward design or introduce extra reward models, which is not the main emphasis of our paper.
We adopt a group size of 12 for both the GRPO and our proposed KRPO models. In our implementation, the Kalman filter state is reset for each prompt to avoid cross-prompt information leakage and ensure statistical independence across prompts. The training batch size is set to 16. The model is optimized using the Adam optimizer with an initial learning rate of $5\times 10^{-6}$. To make the computation manageable, we adopt Llama-3.2-1B-Instruct, Qwen2.5-0.5B-Instruct and Qwen2.5-1.5B-Instruct as our base models, to individually train with the GRPO algorithm and the proposed KRPO algorithm. By default, the Kullback–Leibler (KL) divergence term is weighted by a factor $\alpha$ of 0.01 during training; variance of the filter process noise $Q$ is set to 1e-5 and the variance of the filter measurement noise $R$ is set to 1e-2. For the Arithmetic dataset, the model is trained for 16,000 steps. For the OpenMathInstruct dataset, training is conducted for 5,500 steps. For both MATH500 and AIME datasets, we train the model for 10,000 steps. For the reward design, following the DeepSeek-R1-Zero model (Guo et al., 2025) to keep simplicity, we do not use a reward model and instead assign rewards as follows: if an answer is exactly equal to the ground-truth answer, the reward is 1.0; if a ground-truth answer is contained within a predicted answer, a 0.5 reward is assigned; otherwise, no reward will be returned. We also encourage the model to think reasonably through “<think></think>” tags. The random seed of experiments is set to 42. All experiments are conducted on one A6000 Nvidia GPU with 48GB of graphics memory. The proposed KRPO model and its compared models are trained and tested under the same setting to keep fair comparisons.
4.2 Experimental Results
4.2.1 Overall Model Performance
Table 1: Model accuracy comparison between the KRPO and other models on the Arithmetic and OpenMath-Instruct datasets. The best performance for each block with the same setting is bolded.
@@ 4.2.1 Overall Model Performance
We also show the statistical significance (p-value) of performance gaps between each model and the proposed KRPO model. All differences are statistically significant with p-values below the threshold 0.05, indicating that the observed improvements of KRPO to other models are large. These results collectively indicate that incorporating a Kalman Filter model for advantage estimation is especially beneficial in high-variance, complex reasoning environments. This aligns with the goal of improving the stability of the advantage computation during training. More performance comparisons with other models, including Direct Preference Optimization (DPO) (Rafailov et al., 2023) and fixed baseline models, can be found in the appendix.
4.2.2 Training Reward Curve Comparison between KRPO and GRPO
@@
Additionally, we conduct experiments on additional AIME and AMC data (with Llama3.2-1B-Instruct as the base model) as shown in Fig. 3(b). Despite these two datasets being much more challenging for the Llama3.2-1B model than Arithmetic and OpenMath-Instruct, the results showed similar relative comparisons between KRPO and GRPO under the exactly same conditions: KRPO outperforms GRPO baseline consistently over training steps. More detailed quantitative model performance comparisons on different base models and datasets can be found in the appendix.
@@ 4.3 Analysis
In this sub-section, we analyze the effect of different hyper-parameters and observed phenomena (time efficiency analysis can be found in the appendix). For all experiments in the analysis section are conducted on Arithmetic data.
@@ 4.3.3 Case Study
5 Related Work
5.1 Deep Reinforcement Learning
Combining both policy (Mnih et al., 2013; Van Hasselt et al., 2016; Wang et al., 2016) and value (Sutton et al., 1999; Schulman et al., 2015; 2017) learning, actor-critic methods were proposed. This family includes Advantage Actor-Critic (A2C), which uses a critic to estimate the advantage function, and its asynchronous version Asynchronous Advantage Actor-Critic (A3C) (Mnih et al., 2016), which improves stability and exploration with multi-thread training. PPO can also be viewed as an actor-critic method with a clipped objective. In the continuous action setting, the Deep Deterministic Policy Gradient (DDPG) (Lillicrap et al., 2015) learns a deterministic actor and a critic. Twin Delayed Deep Deterministic Policy Gradient (TD3) (Fujimoto et al., 2018) improves DDPG by addressing function overestimation and policy update delays. Soft Actor-Critic (SAC) (Haarnoja et al., 2018) introduces entropy regularization to encourage exploration and improve stability.
These Reinforcement Learning algorithms can perform well for action control, e.g., playing Atari games (Van Hasselt et al., 2016), MoJoCo (Todorov et al., 2012), or other different simulations (Wang et al., 2021; 2020). However, they are not specifically designed for large language models (LLMs). Reducing the variance of policy gradient estimates is important for achieving stable and sample-efficient training. One widely used method is to subtract a baseline from the observed returns. This baseline typically represents the expected return over all possible actions given the current state. With this subtraction, the resulting action advantage estimate should align more closely with the expected return under the current policy. In commonly used algorithms like Actor-Critic and Proximal Policy Optimization, this baseline is usually obtained from an auxiliary value function model. Nevertheless, maintaining and training this additional model significantly increases computation, particularly when working with large-scale models.
5.2 Reinforcement Learning to Improve Reasoning for LLMs
However, RLHF presents challenges, including the complexity of reward model training and the instability of reinforcement learning procedures. Alternative methods have been proposed to address these issues. Direct Preference Optimization (DPO) (Rafailov et al., 2023) eliminates the need for an explicit reward model by directly optimizing the policy to align with human preferences using a classification loss. This simplification leads to more stable and efficient training. Kahneman-Tversky Optimization (KTO) (Ethayarajh et al., 2024) incorporates principles from prospect theory to better model human decision-making biases. By aligning the training objective with human utility functions, KTO aims to improve the alignment of LLM outputs with human preferences. Reinforced Self-Training (ReST) (Gulcehre et al., 2023) adopts an offline reinforcement learning approach, generating training data from the model’s own outputs and refining the policy without the need for online interaction or extensive human feedback. Rank Responses with Human Feedback (RRHF) (Yuan et al., 2023) simplifies the alignment process by ranking multiple model-generated responses and optimizing the model to prefer higher-ranked outputs, reducing the dependency on complex reward models. Reinforcement Learning from AI Feedback (RLAIF) (Lee et al., 2023) replaces human feedback with evaluations from a separate LLM, enabling scalable alignment without the need for human-labeled data. Self-Play Fine-Tuning (SPIN) (Chen et al., 2024) leverages self-play mechanisms, allowing the model to generate and learn from its own data, progressively improving performance without additional human supervision. Group Relative Policy Optimization (GRPO) (Shao et al., 2024) introduces a group-based approach to policy optimization, where the model’s outputs are evaluated relative to a group baseline.
@@ 6 Conclusion
In this work, we improve the model performance over GRPO by introducing KRPO, a Kalman filter-based method for accurate advantage estimation. Our approach replaces the static mean with a Kalman-filtered baseline and uncertainty that adapt to latent reward dynamics. KRPO maintains the simplicity of GRPO and requires no modification to the policy network or additional learning parameters. Empirical results show that KRPO improves training convergence and policy performance across environments with dynamic and noisy rewards. Besides math reasoning tasks, we are also interested in evaluating KRPO on other tasks such as creative writing and summarization. We consider this to be an important direction for future work.
@@ Appendix D More Comparisons with Other Models
Appendix E Time Efficiency

### v4

# Kalman Filter Enhanced Group Relative Policy Optimization for Language Model Reasoning

@@ Abstract
The advantage function is a central concept in RL that helps reduce variance in policy gradient estimates. Recently, for language modeling, Group Relative Policy Optimization (GRPO) was proposed to compute the advantage for each output by subtracting the mean reward, as the baseline, for all outputs in the group. However, it can lead to high variance when the reward advantage is inaccurately estimated. In this work, we propose Kalman Filter Enhanced Group Relative Policy Optimization (KRPO) model, by using lightweight Kalman filtering to dynamically estimate the latent reward baseline and uncertainty. This filtering technique replaces the naive group mean, enabling more adaptive advantage normalization. Our method does not require additional learned parameters over GRPO. This approach offers a simple yet effective way to incorporate group-level uncertainty for advantage estimation, improving policy optimization in settings where highly dynamic reward signals are difficult to model for language models. Through the accuracies and rewards obtained from math question answering and reasoning, we show that using a more adaptive advantage estimation model, KRPO can improve the performance and show more stable return curves upon GRPO. The code is available at https://anonymous.4open.science/r/KRPO-E1D3.
Machine Learning, ICML
@@ 1 Introduction
2 Related Work
2.1 Deep Reinforcement Learning
Combining both policy (Mnih et al., 2013; Van Hasselt et al., 2016; Wang et al., 2016) and value (Sutton et al., 1999; Schulman et al., 2015, 2017) learning, actor-critic methods were proposed. This family includes Advantage Actor-Critic (A2C), which uses a critic to estimate the advantage function, and its asynchronous version Asynchronous Advantage Actor-Critic (A3C) (Mnih et al., 2016), which improves stability and exploration with multi-thread training. PPO can also be viewed as an actor-critic method with a clipped objective. In the continuous action setting, the Deep Deterministic Policy Gradient (DDPG) (Lillicrap et al., 2015) learns a deterministic actor and a critic. Twin Delayed Deep Deterministic Policy Gradient (TD3) (Fujimoto et al., 2018) improves DDPG by addressing function overestimation and policy update delays. Soft Actor-Critic (SAC) (Haarnoja et al., 2018) introduces entropy regularization to encourage exploration and improve stability.
These Reinforcement Learning algorithms can perform well for action control, e.g., playing Atari games (Van Hasselt et al., 2016), MoJoCo (Todorov et al., 2012), or other different simulations (Wang et al., 2021, 2020). However, they are not specifically designed for large language models (LLMs). Reducing the variance of policy gradient estimates is important for achieving stable and sample-efficient training. One widely used method is to subtract a baseline from the observed returns. This baseline typically represents the expected return over all possible actions given the current state. With this subtraction, the resulting action advantage estimate should align more closely with the expected return under the current policy. In commonly used algorithms like Actor-Critic and Proximal Policy Optimization, this baseline is usually obtained from an auxiliary value function model. Nevertheless, maintaining and training this additional model significantly increases computation, particularly when working with large-scale models.
2.2 Reinforcement Learning to Improve Reasoning for LLMs
However, RLHF presents challenges, including the complexity of reward model training and the instability of reinforcement learning procedures. Alternative methods have been proposed to address these issues. Direct Preference Optimization (DPO) (Rafailov et al., 2023) eliminates the need for an explicit reward model by directly optimizing the policy to align with human preferences using a classification loss. This simplification leads to more stable and efficient training. Kahneman-Tversky Optimization (KTO) (Ethayarajh et al., 2024) incorporates principles from prospect theory to better model human decision-making biases. By aligning the training objective with human utility functions, KTO aims to improve the alignment of LLM outputs with human preferences. Reinforced Self-Training (ReST) (Gulcehre et al., 2023) adopts an offline reinforcement learning approach, generating training data from the model’s own outputs and refining the policy without the need for online interaction or extensive human feedback. Rank Responses with Human Feedback (RRHF) (Yuan et al., 2023) simplifies the alignment process by ranking multiple model-generated responses and optimizing the model to prefer higher-ranked outputs, reducing the dependency on complex reward models. Reinforcement Learning from AI Feedback (RLAIF) (Lee et al., 2023) replaces human feedback with evaluations from a separate LLM, enabling scalable alignment without the need for human-labeled data. Self-Play Fine-Tuning (SPIN) (Chen et al., 2024) leverages self-play mechanisms, allowing the model to generate and learn from its own data, progressively improving performance without additional human supervision. Group Relative Policy Optimization (GRPO) (Shao et al., 2024) introduces a group-based approach to policy optimization, where the model’s outputs are evaluated relative to a group baseline. Recent studies also have critically examined R1-Zero-style reinforcement learning (Liu et al., 2025).
@@ 3.2 Kalman-Filtered Advantage Estimation
We propose replacing the naive group mean baseline with a Kalman filter that maintains an online estimate of the expected reward and its variance to obtain a more accurate estimate of the advantages. Each observed return $r_{i}$ is treated as a noisy measurement of the latent expected reward signal $\hat{x}_{i}$, which we cannot obtain directly. The subscript $i$ refers to the index of the i-th sampled observation.
@@ 3.3 KRPO Algorithm
The Kalman Filtering procedure for advantage estimation is fully differentiable and can be implemented efficiently upon the GRPO frameworks. After getting the group rewards, instead of receiving the mean value from the group, we fetch the Kalman filtered baseline and estimated uncertainty $P_{i|i}$. By taking into account the posterior variance estimation $P_{i|i}$, the advantage can be more adaptively modeled.
@@
Hyperparameters: KL weight $\alpha$, group size $n$, batch size $B$, learning rate $\eta$, clipping $\epsilon$, process noise $Q$, measurement noise $R$ Initialize replay buffer $\mathcal{B}$, optimizer $\text{Adam}(\theta,\eta)$ ;
@@
     Use a Kalman Filter to estimate reward baseline $\hat{x}_{i|i}$ and variance $P_{i|i}$ from $\{r_{i}\}_{i=1}^{n}$ ;
@@
Worth noting, in the model training, the observations are completely randomly sampled in each iteration. Moreover, the order of samples in our implementation is arbitrary (e.g., they can be shuffled before processing), so no unintended temporal dependence is imposed on the data. It simply provides a recursive estimator with an accompanying measure of uncertainty. Although this may introduce a small bias, in practice we found the method insensitive to the permutation.
Besides, each Kalman filter operation is applied to groups of rollouts. In the KRPO setting, although the reward values for individual samples are discrete, the grouped reward observations, which aggregate multiple rollouts, are not sparse and are much more continuous in distribution. According to the Central Limit Theorem (2), the sum of i.i.d. sampled rewards (with finite mean and finite variance) tends to follow a Gaussian distribution. This is consistent with the assumptions of our KRPO setting.
Table 1: Model accuracy comparison between the KRPO and other models on the Arithmetic and OpenMath-Instruct datasets. The best performance for each block with the same setting is bolded. The p-value (last column of table1) is calculated from the answer accuracies by one-tailed paired t-tests between each method against KRPO.
@@
4 Connections of Group-mean and Kalman Filtering
$\bar{r}=\frac{1}{n}\sum_{i=1}^{n}r_{i}.$ (7)
5 Experiments
5.1 Implementation Details
We evaluate our method on four math reasoning datasets: Arithmetic (Open-Thought, 2024), OpenMathInstruct (Toshniwal et al., 2024), MATH500 (Lightman et al., 2023), and AIME (Veeraboina, 2024). These datasets cover a wide spectrum of mathematical tasks, ranging from elementary arithmetic to advanced competition-level problems. Following prior work, we apply a standard 80%/20% train–test split and filter overly long inputs when applicable. To facilitate fine-grained analysis, we further organize Arithmetic and OpenMathInstruct problems into easy, normal, and hard subsets based on either expression complexity or annotated difficulty levels. Full dataset details and preprocessing procedures are provided in Appendix A.
Please note in the paper, we focus on math tasks, as math tasks are verifiable with well-defined answers. These tasks are very suitable for verify the effectiveness of KRPO also because they require the model to have adequate reasoning ability for resolving difficult math questions (even not easy for humans), and language model reasoning is the main focus of our KRPO paper. Reward function or reward model designing is not the main emphasis of our paper.
We adopt a group size of 12 for both the GRPO and our proposed KRPO models. In our implementation, the Kalman filter state is reset for each prompt to avoid cross-prompt information leakage and ensure statistical independence across prompts. The training batch size is set to 16. The model is optimized using the Adam optimizer with an initial learning rate of $5\times 10^{-6}$. To make the computation manageable, we adopt Llama-3.2-1B-Instruct, Llama3.2-3B-Instruct, Qwen2.5-0.5B-Instruct and Qwen2.5-1.5B-Instruct, Qwen2.5-7B-Instruct (experimental results w.r.t. Llama3.2-3B-Instruct and Qwen2.5-7B-Instruct can be found in the Appendix B) as our base models, to individually train with the GRPO algorithm and the proposed KRPO algorithm. By default, the Kullback–Leibler (KL) divergence term is weighted by a factor $\alpha$ of 0.01 during training; variance of the filter process noise $Q$ is set to 1e-5 and the variance of the filter measurement noise $R$ is set to 1e-2. For the Arithmetic dataset, the model is trained for 16,000 steps. For the OpenMathInstruct dataset, training is conducted for 5,500 steps. For both MATH500 and AIME datasets, we train the model for 10,000 steps. For the reward design, following the DeepSeek-R1-Zero model (Guo et al., 2025) to keep simplicity, we do not use a reward model and instead assign rewards as follows: if an answer is exactly equal to the ground-truth answer, the reward is 1.0; if a ground-truth answer is contained within a predicted answer, a 0.5 reward is assigned; otherwise, no reward will be returned. We also encourage the model to think reasonably through “<think></think>” tags. The random seed of experiments is set to 42. All experiments are conducted on one A6000 Nvidia GPU with 48GB of graphics memory. The proposed KRPO model and its compared models are trained and tested under the same setting to keep fair comparisons.
@@ 5.1 Implementation Details
5.2 Experimental Results
5.2.1 Overall Model Performance
We also show the statistical significance (p-value) of performance gaps between each model and the proposed KRPO model. All differences are statistically significant with p-values below the threshold 0.05, indicating that the observed improvements of KRPO to other models are large. These results collectively indicate that incorporating a Kalman Filter model for advantage estimation is especially beneficial in high-variance, complex reasoning environments. This aligns with the goal of improving the stability of the advantage computation during training. More performance comparisons with other models, including Direct Preference Optimization (DPO) (Rafailov et al., 2023) and fixed baseline models, can be found in the Appendix D.
@@ 5.2.1 Overall Model Performance
5.2.2 Training Reward Curve Comparison between KRPO and GRPO
@@
Additionally, we conduct experiments on additional AIME and AMC data (with Llama3.2-1B-Instruct as the base model) as shown in Fig. 3(b). Despite these two datasets being much more challenging for the Llama3.2-1B model than Arithmetic and OpenMath-Instruct, the results showed similar relative comparisons between KRPO and GRPO under the exactly same conditions: KRPO outperforms GRPO baseline consistently over training steps. More detailed quantitative model performance comparisons on different base models and datasets can be found in the Appendix B and  C.
@@ 5.3 Analysis
In this sub-section, we analyze the effect of different hyper-parameters and observed phenomena (time efficiency analysis can be found in the Appendix G). For all experiments in the analysis section are conducted on Arithmetic data.
@@ 6 Conclusion
In this work, we improve the model performance over GRPO by introducing KRPO, a Kalman filter-based method for accurate advantage estimation. Our approach replaces the static mean with a Kalman-filtered baseline and uncertainty that adapt to latent reward dynamics. KRPO maintains the simplicity of GRPO and requires no modification to the policy network or additional learning parameters. Empirical results show that KRPO improves training convergence and policy performance across environments with dynamic and noisy rewards.
More broadly, KRPO demonstrates that principled uncertainty-aware estimation can be incorporated into policy optimization with minimal overhead, offering a practical alternative to estimate advantages. In future work, we plan to evaluate KRPO on tasks such as creative writing and summarization, where reward signals are inherently noisier and more subjective.
@@ Appendix B Model Performance on More Base-models
To evaluate the generalization with larger models, we also conducted experiments with Llama3.2-3B-Instruct model as shown in the Fig. 5 and Fig. 6. With the Llama3.2-3B-Instruct and Qwen2.5-7B-Instruct models, the training returns also demonstrate that our proposed KRPO method consistently achieves higher rewards than the GRPO baseline, confirming the effectiveness of our approach.
Figure 5: The training return curves of Llama3.2-3B-Instruct model on normal level of Arithmetic data.
Figure 6: The training return curves of Qwen2.5-7B-Instruct model on normal level of Arithmetic data.
@@ Appendix D More Comparisons with Other Models
Appendix E Model Performance with Different Group Size
An interesting point to be noted: the rollout size may influence the variance reduction behavior of both GRPO and KRPO. Nevertheless, this effect is not unique to KRPO. In GRPO, the group mean baseline also becomes less accurate when the number of rollouts per prompt is small, which is a natural consequence of the Law of Large Numbers via Monte Carlo sampling. The main difference being that KRPO provides an adaptive and stable advantage estimation within each group.
Therefore, we also conducted experiments on model performance with different group sizes (= {5, 8, 10, 15}) as shown in Fig. 7(a). In the figure, “gx” denotes with group size x (e.g. g5 means with group size 5). Across various group size settings, the proposed KRPO consistently outperformed the baseline GRPO, demonstrating its robustness and effectiveness. Interestingly, as the group size varies from 5 to 15, the performance of both KRPO and GRPO improves; however, a large performance gap always remains between them, with KRPO maintaining a clear advantage. Interestingly, we empirically observe that as the number of group samples are very few, the performance gap between KRPO and its GRPO baseline widens, indicating that KRPO is particularly effective at handling extremely noisy rewards when only very sparse group samples are available.
Appendix F Model Performance with Different Random Seeds
(a) Model performance with different group sizes.
(b) Model performance with different random seeds.
Figure 7: The training return curves of different (a) group sizes and (b) seeds. The base model is Llama-3.2-1B-Instruct on normal level of Arithmetic data.
We conducted experiments under multiple random seed settings to verify the robustness of the results, as shown in Fig. 7(b). In the figure, “sx” presents with seed x (e.g. s777 means with random seed 777). We observe that when the random seeds are set to 777 and 1234, similar to the default seed 42, KRPO yields consistently higher collected rewards than GRPO, indicating that the superiority of KRPO is not sensitive to the choice of random initialization.
Appendix G Time Efficiency